# RoleInteract: Evaluating the Social Interaction of Role-Playing Agents

### v3

@@ Abstract
Large language models (LLMs) have advanced the development of various AI conversational agents, including role-playing conversational agents that mimic diverse characters and human behaviors. While prior research has predominantly focused on enhancing the conversational capability, role-specific knowledge, and stylistic attributes of these agents, there has been a noticeable gap in assessing their social intelligence. In this paper, we introduce RoleInteract, the first benchmark designed to systematically evaluate the sociality of role-playing conversational agents at both individual and group levels of social interactions. The benchmark is constructed from a variety of sources and covers a wide range of 500 characters and over 6,000 question prompts and 30,800 multi-turn role-playing utterances. We conduct comprehensive evaluations on this benchmark using mainstream open-source and closed-source LLMs. We find that agents excelling in individual level does not imply their proficiency in group level. Moreover, the behavior of individuals may drift as a result of the influence exerted by other agents within the group. Experimental results on RoleInteract confirm its significance as a testbed for assessing the social interaction of role-playing conversational agents. The benchmark is publicly accessible at https://github.com/X-PLUG/RoleInteract.
RoleInteract: Evaluating the Social Interaction of Role-Playing Agents
@@ 1 Introduction
Therefore, we introduce RoleInteract, the first evaluation benchmark designed to systematically assess the social interaction of role-playing conversational agents. As introduced in Troitzsch (1996), the agent society represents a complex system comprising individual and group social activities. Following this definition, RoleInteract assesses the sociality metrics at both the individual and group levels, as shown in Figure 1. At the individual level, the agent should possess the basic social intelligence as individuals, such as self-awareness on role description (Wang et al., 2023c; Tu et al., 2024; Shen et al., 2023), emotional perception on environment (Hsu et al., 2018), and long-term conversation memory (Zhong et al., 2023). Each of these aspects contributes to the nuanced understanding of how the agents manifest their individual social behaviors. Moreover, we further examine the social intelligence of the role-playing agents within group social interactions, which require the agents to possess certain social preferences towards group dynamics (Leng et al., 2023).
RoleInteract is carefully constructed from diverse English and Chinese books, movies, and novels, covering a wide range of 500 characters and 6,000 questions, and 30,800 multi-turn role-playing utterances. Specifically, we design a three-step construction pipeline for RoleInteract. Firstly, we collect diverse role profiles from common web sources. Secondly, GPT-4 is employed to extract dialogue scenes, individual and group-level social conversations, as well as multi-choice questions. Thirdly, we conduct a series of pre-processing and manual labeling to ensure the quality of the benchmark. we conduct comprehensive evaluations on RoleInteract using mainstream open-source and closed-source LLMs to inspire future research.
@@ 2.2 Role-Playing Benchmarks
With the rapid development of role-playing agents, there has been a corresponding increase in evaluation datasets. Current evaluation datasets mostly focus on the alignment of role-playing agents with regards to role style and role knowledge. In terms of role style, Tu et al. (2024) and Wang et al. (2023c) investigate whether models can generate responses consistent with the style of the given role. Agents need to grasp different speaking styles for different roles. Regarding role knowledge, Shen et al. (2023) particularly focuses on the role knowledge of role-playing models, including the characters’ experiences and social relationships. Tu et al. (2024) and Wang et al. (2023c) also address aspects of role knowledge, such as role knowledge illusions. Additionally, Wang et al. (2023a) and Tu et al. (2024) introduce psychological theories like the Big Five and MBTI to evaluate role-playing agents. While previous work mainly focuses on testing the abilities of agents on imitating the character’s role-specific knowledge or speaking style, RoleInteract introduces the first-ever evaluation benchmark for the sociality of role-playing conversational agents encompassing both individual and group level.
@@ 2.3 Agent Society
Previous benchmarks have primarily focused on single-agent scenarios, leaving the more complex multi-agent scenarios underexplored. Similar to humans, agents are capable of engaging in intricate social interactions, resulting in the formation of an agent society (da Rocha Costa, 2019). Recently, LLM-based agents demonstrate complex social behaviors, where cooperation and competition coexist (Xu et al., 2023). These sophisticated behaviors intertwine to shape social interactions (Gao et al., 2023). RoleInteract follows the framework defined by Nigel Gilbert and Troitzsch (1997); Leng et al. (2023), where behaviors in agent societies are divided into individual and group-level activities, to study the social intelligence of role-playing agents within social interactions.
Figure 1: An example from RoleInteract, which is partially constructed from the film “The Great Gatsby”.
@@ 3.2 Group Level
4 RoleInteract
In this section, we introduce the construction process of RoleInteract, as illustrated in Figure 2.
@@ 4.1 Profile Collection
A role profile defines the character style, knowledge, emotions, and social preference of a role-playing agent. We gather profiles for role-playing agents from various sources including novels, scripts, online platforms such as CharacterAI [^2] https://beta.character.ai and Fandom [^3] https://www.fandom.com, and automatic generation via GPT-4-Turbo prompting. To ensure diversity, we construct profiles based on various character types and personality traits by combining the existing categorizations in online platforms and research work (Shen et al., 2023; Gunkel, 1998). Figure 3 illustrates the distribution of personality traits for roles within RoleInteract. It demonstrates our approach of ensuring both category diversity and balanced proportions across all categories. The details regarding profile collection in RoleInteract can be found in Appendix D.1.
Figure 2: The three-step dataset construction pipeline of RoleInteract.
Figure 3: Personality traits distribution in RoleInteract.
@@ 4.3 Question Design
Based on the constructed dialogues, we employ different methods for designing questions tailored to different dimensions within RoleInteract.
@@ 4.4.2 Dataset Post-Validation
Table 1: Metrics and statistics of RoleInteract. There are a total of 500 roles, comprising 6,000 questions and 30,800 utterances in RoleInteract.
@@ 5 Experiment Settings
In this section, we show the statistic of RoleInteract. Then we introduce the metrics along with the evaluation LLMs.
@@ 5.1 Dataset Statistic
We show the statistic of RoleInteract in Table 1 and the distribution of dialogue tokens length in Figure 4. RoleInteract encompasses individual level and group level. There are six subcategories in individual level: self-awareness on role style (SA Style), self-awareness on role knowledge (SA Know), situational understanding (EP Situ.), emotion detection (EP Emo.), short-term conversation memory (CM Short), and long-term conversation memory (CM Long). Group level consists of three social preference categories: positive (Pos.), neutral (Neu.), and negative (Neg.). There are a total of 500 roles, comprising 6,000 questions and 30,800 utterances in RoleInteract.
@@ 5.2 Evaluation Metrics
Most of the previous methods (Wang et al., 2023c; Shao et al., 2023) for role-playing applications rely on GPT-3.5 or GPT-4 for evaluation, which may suffer from questionable accuracy on the role-playing scenario and costly API usage. We follow the popular benchmark MMLU Hendrycks et al. (2020) and C-Eval Huang et al. (2023), and prompt for automatic and fast evaluation free from LLMs. RoleInteract utilizes fully automatic evaluation metrics, employing both multiple-choice and open-domain generation questions.
@@
For open-domain question, we calculate the keyword coverage rate ($𝑜𝑣𝑒𝑟C$). RoleInteract provides a keyword set $\mathbf{A}_{\text{keywords}}=\{k_{1},k_{2},\dots,k_{n}\}$. Given the keywords set mentioned in the response $\mathbf{R}_{\text{keywords}}$, we compute:
@@
The metrics utilized for different dimensions in RoleInteract are listed in Table 1.
Figure 4: Distribution of dialogue tokens across four dimensions in RoleInteract, based on tokenizer of Qwen.
@@
Table 2: Main results from RoleInteract. Best performances are shown in bold, while suboptimal ones underlined.
@@ 6.2 Conversation Memory for Role-Playing
Conversation memory capability is crucial for role-playing agents. We investigate the memory capacity of role-playing agents across different conversation lengths, measured by the number of utterances in the dialogue. We analyze the distribution of utterance counts in the conversation memory dimension of RoleInteract. As illustrated in Figure 5, there is a declining trend in memory capability for some models, such as GPT-3.5-Turbo and CharGLM-3, as conversation length increases. When the number of utterances in the dialogue exceeds 80 rounds, most role-playing agents exhibit a noticeable performance decline. This finding showcases the limitations of current role-playing agents in handling extremely long-term memory and highlights potential areas for improvement.
@@ 6.4 Impact of Group Dynamics Polarity
It is important for role-playing agents to maintain designed social preferences under the influence of varying group dynamics. The group dynamics polarity is defined as the majority social preference of group members. For instance, positive group dynamics imply that the majority of members exhibit positive social preference. For an individual with a specific social preference, different polarities of group dynamics may have various impacts. We study the performance of individuals under different polarities of group dynamics, by analyzing a subset of group data in RoleInteract. As shown in Figure 7, we find that individuals with neutral and negative social preferences perform optimally within their corresponding group polarities (i.e., neutral and negative group polarities). However, they are susceptible to the influence of group dynamics with different polarities and undergo a phenomenon termed as preference drift, leading to deviation from their original designed behaviors, as indicated by the decline of performance. Nevertheless, individuals with positive social preference appear to be more resilient to the preference drift, performing better across all group polarities. Especially, they excel in negative group polarity. This phenomenon can be termed as social facilitation (Guerin, 2010) in sociology. We hypothesize that negative group further motivates individuals to engage in behaviors advantageous to the group.
@@ 7 Conclusion
In this paper, we introduce RoleInteract, the first evaluation benchmark designed to systematically assess the social intelligence of role-playing conversational agents at both individual and group levels. We construct diverse question prompts on a wide range of characters covering comprehensive dimensions, including self-awareness on role description, emotional perception on environment, long-term conversation memory, and social preference towards group dynamics. Moreover, rigorous human verification ensure questions’ difficulty and validity. We evaluate over 10 mainstream LLMs on RoleInteract and provide in-depth analysis. While role-playing agents demonstrate satisfactory performance at the individual level, we find that their social interaction capabilities at the group level remain deficient. We hope this finding may inspire future research in this field.
@@
While RoleInteract provides a comprehensive evaluation framework for assessing the sociality of role-playing conversation agents, there are several limitations to consider. 1) Social interactions, particularly within group settings, are inherently complex and nuanced. Despite our efforts, further research is needed to fully understand and capture the intricacies of these interactions. 2) The number of role-playing agents in group scenarios is relatively limited in our benchmark. Increasing the diversity and quantity of agents would provide a more comprehensive evaluation of the agents’ social abilities and dynamics within groups. 3) Our dataset may contain some biased content, posing a risk of improper use. These limitations highlight areas for future research and development in the evaluation of social intelligence in role-playing agents.
Appendix A Examples from RoleInteract
We showcase examples from RoleInteract in Figures 8, 9, 10, and 11. A typical example consists of a character’s profile, conversation history, instruction, and question. There may be differences in format across certain dimensions. For example, in the emotional perception dimension, there is no character profile provided. In the conversation memory dimension, answers to each question are in the form of keywords rather than multiple-choice options. The conversation is stored in the format of a list combined with dictionaries. Each utterance is represented as a dictionary, where the keys are the names of the characters and the values are the content spoken by each character.
Figure 8: Example from self-awareness on role style in RoleInteract.
Figure 9: Example from conversation memory dimension in RoleInteract.
Figure 10: Example from social preference dimension in RoleInteract.
Figure 11: Example from emotional perception dimension in RoleInteract.
@@ B.2 Question Design
For conversation memory: This category includes two subcategories: short-term conversation memory (CM Short) and long-term conversation memory (CM Long). In RoleInteract, questions for other dimensions are presented in multiple-choice format. However, to increase the difficulty of the conversation memory dimension, we utilize an open-domain generation combined with keyword matching approach for this dimension. For example, if an agent previously answered that they had a sandwich for breakfast, after several rounds of conversation, if the user asks again what the agent had for breakfast, we require the agent’s response to include the keyword "sandwich." If the agent responds that they had bread for breakfast, since the keyword does not match, we consider the agent unable to correctly recall their previous dialogue content. For CM Short, we prompt the agent to recall keywords discussed within 40 utterances, while for CM Long, we prompt the agent to recall keywords discussed over 40 utterances. We evaluate how many of these keywords are recalled.
@@ B.3 Dataset Validation
Table 7: Personality traits in RoleInteract.
@@ Appendix D Dataset Statistic
RoleInteract covers 500 characters and 6,000 question prompts involved in 1,000 conversation scenarios and 30,800 multi-turn role-playing utterances.
@@ D.1 Personality Traits
We follow the definition of personality traits in Gunkel (1998) to construct profiles, ensuring diversity and comprehensiveness in RoleInteract. From the collection of 638 personality descriptors created by Gunkel (1998), we selected a subset of easily understandable terms for construction. These selected terms can be categorized into positive, neutral, and negative traits, as illustrated in Table 7.

### v4

# SocialBench: Sociality Evaluation of Role-Playing Conversational Agents

@@ Abstract
Large language models (LLMs) have advanced the development of various AI conversational agents, including role-playing conversational agents that mimic diverse characters and human behaviors. While prior research has predominantly focused on enhancing the conversational capability, role-specific knowledge, and stylistic attributes of these agents, there has been a noticeable gap in assessing their social intelligence. In this paper, we introduce SocialBench, the first benchmark designed to systematically evaluate the sociality of role-playing conversational agents at both individual and group levels of social interactions. The benchmark is constructed from a variety of sources and covers a wide range of 500 characters and over 6,000 question prompts and 30,800 multi-turn role-playing utterances. We conduct comprehensive evaluations on this benchmark using mainstream open-source and closed-source LLMs. We find that agents excelling in individual level does not imply their proficiency in group level. Moreover, the behavior of individuals may drift as a result of the influence exerted by other agents within the group. Experimental results on SocialBench confirm its significance as a testbed for assessing the social interaction of role-playing conversational agents. The benchmark is publicly accessible at https://github.com/X-PLUG/SocialBench.
SocialBench: Sociality Evaluation of Role-Playing Conversational Agents
@@ 1 Introduction
Therefore, we introduce SocialBench, the first evaluation benchmark designed to systematically assess the social interaction of role-playing conversational agents. As introduced in Troitzsch (1996), the agent society represents a complex system comprising individual and group social activities. Following this definition, SocialBench assesses the sociality metrics at both the individual and group levels, as shown in Figure 1. At the individual level, the agent should possess the basic social intelligence as individuals, such as self-awareness on role description (Wang et al., 2023c; Tu et al., 2024; Shen et al., 2023), emotional perception on environment (Hsu et al., 2018), and long-term conversation memory (Zhong et al., 2023). Each of these aspects contributes to the nuanced understanding of how the agents manifest their individual social behaviors. Moreover, we further examine the social intelligence of the role-playing agents within group social interactions, which require the agents to possess certain social preferences towards group dynamics (Leng et al., 2023).
SocialBench is carefully constructed from diverse English and Chinese books, movies, and novels, covering a wide range of 500 characters and 6,000 questions, and 30,800 multi-turn role-playing utterances. Specifically, we design a three-step construction pipeline for SocialBench. Firstly, we collect diverse role profiles from common web sources. Secondly, GPT-4 is employed to extract dialogue scenes, individual and group-level social conversations, as well as multi-choice questions. Thirdly, we conduct a series of pre-processing and manual labeling to ensure the quality of the benchmark. we conduct comprehensive evaluations on SocialBench using mainstream open-source and closed-source LLMs to inspire future research.
@@ 2.2 Role-Playing Benchmarks
With the rapid development of role-playing agents, there has been a corresponding increase in evaluation datasets. Current evaluation datasets mostly focus on the alignment of role-playing agents with regards to role style and role knowledge. In terms of role style, Tu et al. (2024) and Wang et al. (2023c) investigate whether models can generate responses consistent with the style of the given role. Agents need to grasp different speaking styles for different roles. Regarding role knowledge, Shen et al. (2023) particularly focuses on the role knowledge of role-playing models, including the characters’ experiences and social relationships. Tu et al. (2024) and Wang et al. (2023c) also address aspects of role knowledge, such as role knowledge illusions. Additionally, Wang et al. (2023a) and Tu et al. (2024) introduce psychological theories like the Big Five and MBTI to evaluate role-playing agents. While previous work mainly focuses on testing the abilities of agents on imitating the character’s role-specific knowledge or speaking style, SocialBench introduces the first-ever evaluation benchmark for the sociality of role-playing conversational agents encompassing both individual and group level.
@@ 2.3 Agent Society
Previous benchmarks have primarily focused on single-agent scenarios, leaving the more complex multi-agent scenarios underexplored. Similar to humans, agents are capable of engaging in intricate social interactions, resulting in the formation of an agent society (da Rocha Costa, 2019). Recently, LLM-based agents demonstrate complex social behaviors, where cooperation and competition coexist (Xu et al., 2023). These sophisticated behaviors intertwine to shape social interactions (Gao et al., 2023). SocialBench follows the framework defined by Nigel Gilbert and Troitzsch (1997); Leng et al. (2023), where behaviors in agent societies are divided into individual and group-level activities, to study the social intelligence of role-playing agents within social interactions.
Figure 1: An example from SocialBench, which is partially constructed from the film “The Great Gatsby”.
@@ 3.2 Group Level
4 SocialBench
In this section, we introduce the construction process of SocialBench, as illustrated in Figure 2.
@@ 4.1 Profile Collection
A role profile defines the character style, knowledge, emotions, and social preference of a role-playing agent. We gather profiles for role-playing agents from various sources including novels, scripts, online platforms such as CharacterAI [^2] https://beta.character.ai and Fandom [^3] https://www.fandom.com, and automatic generation via GPT-4-Turbo prompting. To ensure diversity, we construct profiles based on various character types and personality traits by combining the existing categorizations in online platforms and research work (Shen et al., 2023; Gunkel, 1998). Figure 3 illustrates the distribution of personality traits for roles within SocialBench. It demonstrates our approach of ensuring both category diversity and balanced proportions across all categories. The details regarding profile collection in SocialBench can be found in Appendix D.1.
Figure 2: The three-step dataset construction pipeline of SocialBench.
Figure 3: Personality traits distribution in SocialBench.
@@ 4.3 Question Design
Based on the constructed dialogues, we employ different methods for designing questions tailored to different dimensions within SocialBench.
@@ 4.4.2 Dataset Post-Validation
Table 1: Metrics and statistics of SocialBench. There are a total of 500 roles, comprising 6,000 questions and 30,800 utterances in SocialBench.
@@ 5 Experiment Settings
In this section, we show the statistic of SocialBench. Then we introduce the metrics along with the evaluation LLMs.
@@ 5.1 Dataset Statistic
We show the statistic of SocialBench in Table 1 and the distribution of dialogue tokens length in Figure 4. SocialBench encompasses individual level and group level. There are six subcategories in individual level: self-awareness on role style (SA Style), self-awareness on role knowledge (SA Know), situational understanding (EP Situ.), emotion detection (EP Emo.), short-term conversation memory (CM Short), and long-term conversation memory (CM Long). Group level consists of three social preference categories: positive (Pos.), neutral (Neu.), and negative (Neg.). There are a total of 500 roles, comprising 6,000 questions and 30,800 utterances in SocialBench.
@@ 5.2 Evaluation Metrics
Most of the previous methods (Wang et al., 2023c; Shao et al., 2023) for role-playing applications rely on GPT-3.5 or GPT-4 for evaluation, which may suffer from questionable accuracy on the role-playing scenario and costly API usage. We follow the popular benchmark MMLU Hendrycks et al. (2020) and C-Eval Huang et al. (2023), and prompt for automatic and fast evaluation free from LLMs. SocialBench utilizes fully automatic evaluation metrics, employing both multiple-choice and open-domain generation questions.
@@
For open-domain question, we calculate the keyword coverage rate ($𝑜𝑣𝑒𝑟C$). SocialBench provides a keyword set $\mathbf{A}_{\text{keywords}}=\{k_{1},k_{2},\dots,k_{n}\}$. Given the keywords set mentioned in the response $\mathbf{R}_{\text{keywords}}$, we compute:
@@
The metrics utilized for different dimensions in SocialBench are listed in Table 1.
Figure 4: Distribution of dialogue tokens across four dimensions in SocialBench, based on tokenizer of Qwen.
@@
Table 2: Main results from SocialBench. Best performances are shown in bold, while suboptimal ones underlined.
@@ 6.2 Conversation Memory for Role-Playing
Conversation memory capability is crucial for role-playing agents. We investigate the memory capacity of role-playing agents across different conversation lengths, measured by the number of utterances in the dialogue. We analyze the distribution of utterance counts in the conversation memory dimension of SocialBench. As illustrated in Figure 5, there is a declining trend in memory capability for some models, such as GPT-3.5-Turbo and CharGLM-3, as conversation length increases. When the number of utterances in the dialogue exceeds 80 rounds, most role-playing agents exhibit a noticeable performance decline. This finding showcases the limitations of current role-playing agents in handling extremely long-term memory and highlights potential areas for improvement.
@@ 6.4 Impact of Group Dynamics Polarity
It is important for role-playing agents to maintain designed social preferences under the influence of varying group dynamics. The group dynamics polarity is defined as the majority social preference of group members. For instance, positive group dynamics imply that the majority of members exhibit positive social preference. For an individual with a specific social preference, different polarities of group dynamics may have various impacts. We study the performance of individuals under different polarities of group dynamics, by analyzing a subset of group data in SocialBench. As shown in Figure 7, we find that individuals with neutral and negative social preferences perform optimally within their corresponding group polarities (i.e., neutral and negative group polarities). However, they are susceptible to the influence of group dynamics with different polarities and undergo a phenomenon termed as preference drift, leading to deviation from their original designed behaviors, as indicated by the decline of performance. Nevertheless, individuals with positive social preference appear to be more resilient to the preference drift, performing better across all group polarities. Especially, they excel in negative group polarity. This phenomenon can be termed as social facilitation (Guerin, 2010) in sociology. We hypothesize that negative group further motivates individuals to engage in behaviors advantageous to the group.
@@ 7 Conclusion
In this paper, we introduce SocialBench, the first evaluation benchmark designed to systematically assess the social intelligence of role-playing conversational agents at both individual and group levels. We construct diverse question prompts on a wide range of characters covering comprehensive dimensions, including self-awareness on role description, emotional perception on environment, long-term conversation memory, and social preference towards group dynamics. Moreover, rigorous human verification ensure questions’ difficulty and validity. We evaluate over 10 mainstream LLMs on SocialBench and provide in-depth analysis. While role-playing agents demonstrate satisfactory performance at the individual level, we find that their social interaction capabilities at the group level remain deficient. We hope this finding may inspire future research in this field.
@@
While SocialBench provides a comprehensive evaluation framework for assessing the sociality of role-playing conversation agents, there are several limitations to consider. 1) Social interactions, particularly within group settings, are inherently complex and nuanced. Despite our efforts, further research is needed to fully understand and capture the intricacies of these interactions. 2) The number of role-playing agents in group scenarios is relatively limited in our benchmark. Increasing the diversity and quantity of agents would provide a more comprehensive evaluation of the agents’ social abilities and dynamics within groups. 3) Our dataset may contain some biased content, posing a risk of improper use. These limitations highlight areas for future research and development in the evaluation of social intelligence in role-playing agents.
Appendix A Examples from SocialBench
We showcase examples from SocialBench in Figures 8, 9, 10, and 11. A typical example consists of a character’s profile, conversation history, instruction, and question. There may be differences in format across certain dimensions. For example, in the emotional perception dimension, there is no character profile provided. In the conversation memory dimension, answers to each question are in the form of keywords rather than multiple-choice options. The conversation is stored in the format of a list combined with dictionaries. Each utterance is represented as a dictionary, where the keys are the names of the characters and the values are the content spoken by each character.
Figure 8: Example from self-awareness on role style in SocialBench.
Figure 9: Example from conversation memory dimension in SocialBench.
Figure 10: Example from social preference dimension in SocialBench.
Figure 11: Example from emotional perception dimension in SocialBench.
@@ B.2 Question Design
For conversation memory: This category includes two subcategories: short-term conversation memory (CM Short) and long-term conversation memory (CM Long). In SocialBench, questions for other dimensions are presented in multiple-choice format. However, to increase the difficulty of the conversation memory dimension, we utilize an open-domain generation combined with keyword matching approach for this dimension. For example, if an agent previously answered that they had a sandwich for breakfast, after several rounds of conversation, if the user asks again what the agent had for breakfast, we require the agent’s response to include the keyword "sandwich." If the agent responds that they had bread for breakfast, since the keyword does not match, we consider the agent unable to correctly recall their previous dialogue content. For CM Short, we prompt the agent to recall keywords discussed within 40 utterances, while for CM Long, we prompt the agent to recall keywords discussed over 40 utterances. We evaluate how many of these keywords are recalled.
@@ B.3 Dataset Validation
Table 7: Personality traits in SocialBench.
@@ Appendix D Dataset Statistic
SocialBench covers 500 characters and 6,000 question prompts involved in 1,000 conversation scenarios and 30,800 multi-turn role-playing utterances.
@@ D.1 Personality Traits
We follow the definition of personality traits in Gunkel (1998) to construct profiles, ensuring diversity and comprehensiveness in SocialBench. From the collection of 638 personality descriptors created by Gunkel (1998), we selected a subset of easily understandable terms for construction. These selected terms can be categorized into positive, neutral, and negative traits, as illustrated in Table 7.